OpenStudy (anonymous):

Find the equation of the tangent line to the curve y = 2 x cos x at the point ( pi , -2 \pi). The equation of this tangent line can be written in the form y = mx+b where m = b =

OpenStudy (kira_yamato):

y = 2 x cos x y' = 2cos x - 2x sinx y'(π) = 2cosπ - 2π sinπ = -2 y + 2π = -2 (x - π) y = -2x

OpenStudy (anonymous):

where's m and b?